Calculate number of electrons present in 9.5 g of phosphate( PO^3-{base4} )
1
Expert's answer
2020-11-12T06:35:53-0500

Number of moles,

9.5÷95=0.1

Total electrons 7 +(8×4)+3=50

0.1×50×6×1023=3×1024